26 cc of CO2 is passed over red hot coke. What is the volume of CO evolved out?

Grade:11

2 Answers

Riddhi Jain
7 years ago
Co2 +C=CO. 22.4 l gives 2×22.4 l of co2....................therefore 26cc will give 26×2×22.4/22.4= 52 cc of co2
SAURABH SONI
5 years ago
Co2 + C =2CO
22.4 lit co2 gives 2*22.4 lit of CO
1lit---------2*22.4/22.4
26*10-३  lit----- 22.4*2*26/22.4=52